3/4 as ?/16y-- i can do it .. except they threw in the letter y.. that threw me off.. please help
Answer by uma(370)

3/4 = ?/16y
This is simple.
Divide 16y by 4.
We get 4y.
Then 3 * 4y = 12y
Thus 3/4 = 12y/16y

Good Luck!!!